# Discovery of energy landscapes towards optimized quantum transport: Environmental effects and long-range tunneling

### v1

@@ IV.3 OQS Model II
We continue our investigation of transport, optimizing structures at finite temperatures using OQS Model II, as described by Eq. (6). The results are shown in Figure 10(a)-(b). Unlike the Coherent Model and OQS Model I, we identify here several consistent trends across both short- and long-range tunneling chains. In both cases, the optimal energy profile takes on a ramp-like steadily decreasing form. This configuration can be rationalized: at moderate temperatures, it promotes forward transfer from the entrance to the exit site while effectively suppressing backflow. For short-range tunneling, we also observe the emergence of high-energy levels above both the entrance and exit sites. Similarly to the three-site models shown in Figs. 4(c) and (f), this arrangement reduces population buildup along the chain itself while enhancing accumulation at the exit site.
Figure 10(c)-(d) show the population distributions corresponding to the optimized energy profiles, while Fig. 10(e)-(h) presents the off-diagonal elements of the density matrix. Interestingly, after the first one or two sites, the population in the chain increases towards the exit site, indicating on the accumulation of carriers towards the end of the chain due to the ramp potential. Compared to optimal structures identified in OQS Model I, coherences in the finite-temperature model are more localized, extending only to about three neighboring sites for both short- and long-range tunneling cases. This is consistent with the ramp structure employed in OQS Model II. To highlight the robustness of these findings, Fig. 11 compiles a set of energy profiles achieved during the optimization process, all consistently exhibiting the characteristic ramp-like structure.
@@
Model Energy landscape Population profile Coherence trends Proposed mechanism
Coherent model, $\alpha=1$ nearly flat nearly constant localized ballistic
Coherent model, $\alpha=3$ corrugated bowl shaped extended bridge-assisted tunneling
OQS Model I, $\alpha=1$ nearly flat decaying approximately linearly localized diffusive
OQS Model I, $\alpha=3$ corrugated constant for $n>4$ extended ballistic for long chains
OQS Model II, $\alpha=1$ decreasing ramp growing towards exit site localized hopping
OQS Model II, $\alpha=3$ decreasing ramp about constant at the center localized hopping
Table 1: Summary of models, energy landscapers optimizing transport, and the associated transport mechanism.
@@ V Conclusions
We identified distinct classes of energy landscapes that maximize population flux along quasi-one-dimensional chains. Hand-in-hand with the optimization, the analysis of the resulting steady state allowed us to uncover transport mechanisms. Towards the objective of optimized transport, we specifically compared (i) the role of the environment and (ii) the impact of short-range vs. long-range tunneling on chain design. The interplay of these incoherent and coherent effects leads to complex transport behavior. Optimized transport in the models required these factors to cooperate.
@@
We modeled the chain’s interaction with its environment using Lindblad jump operators. In OQS Model I, we incorporated local Lindblad dephasing, which, when viewed in the energy eigenbasis, corresponds to coupling with an infinite-temperature bath. In contrast, OQS Model II represents a finite-temperature environment that drives both excitation and relaxation processes between energy eigenstates, while satisfying the detailed balance condition.
@@
We identified the following design rules for optimizing transfer, summarized in Table I.
@@
(i) Coherent model: For short-range tunneling, the optimal configuration features an almost uniform energy landscape, enabling near-ballistic transport. When long-range tunneling contributes, the optimal energy profile is non-uniform and corrugated, with intermediate sites energetically detuned. This facilitates a more direct transfer between the chain’s endpoints, as evidenced by strong steady-state coherences linking those sites. Long-range coherent tunneling systems are highly sensitive to variations in parameters near optimal energy landscapes.
@@
(ii) OQS Model I: When local dephasing is significant and comparable in magnitude to the tunneling energy, models with short-range tunneling favor nearly flat energy landscapes, with only slight detuning from the entry and exit sites to facilitate conduction. In contrast, under long-range tunneling, a corrugated energy landscape allows the system to cross over to ballistic-like conduction away from the entrance site by suppressing interferences and facilitating efficient end-to-end transfer.
@@
(iii) OQS Model II: At moderate temperatures ($T\approx J_{max}$), ramp-like energy profiles with a total gap of the order of $J_{max}$ are preferred, regardless of tunneling range, as they support directional transport and improve the collection of carriers at the exit site. In these systems, transport proceed in an almost localized manner with small coherences.
@@
The optimization process posed challenges at times, especially for the coherent model and OQS Model II. In the coherent model, particularly with long-range tunneling, the steady-state populations were highly sensitive to small changes in energy levels, resulting in complex, rugged flux landscapes with numerous local maxima that complicated optimization. For OQS Model II, the need to re-diagonalize the system Hamiltonian at each energy iteration further increased computational demands, making the optimization considerably slower compared to OQS Model I under the same parameters.
@@
Long-range steady-state coherences, such as those found under OQS Model I in Fig. 7(f)-(h) have been shown to lead to robust entanglement [100, 94, 101]. As such, the energy profiles we obtain under long-range tunneling [purple in Fig. 7(a)-(b)] may find applications in quantum information processing and quantum metrology [94, 102, 103].
@@
In future work, we aim to extend our study of carrier transfer to situations involving multiple carriers, such as transport in spin systems modeled by Heisenberg-type chains with long-range exchange interactions. Another promising direction is the investigation of excitation transfer in molecular aggregates or disordered materials coupled to an optical cavity [104, 105, 106, 83]. Such cavity coupling can enable long-range excitation transfer. Joint optimization of both the molecular structure and its interaction with the cavity may yield new design principles for efficient carrier transport in quantum networks.
@@ Appendix A Details of the optimization process
Stochastic Gradient Descent (SGD) and Noisy SGD algorithms were also tested with three sites, the results of which are shown in Figs. 12(b) and 12(d), respectively. Both found the same optima as OGA and AdaMax, but with more steps: SGD took between 890 and 17,700 steps, and with the default noise levels, Noisy SGD took between 19,000 and 28,000 steps.
@@
Figure 12: Examples of optimization trajectories in a three-site system under OQS Model I. Squares mark the initial energies supplied to the optimizer, while stars indicate where the optimizers hit the stopping condition. Default Optax parameters were used for each algorithm, except for the learning rate. The optimizers are: (a) optimistic gradient ascent, learning rate 0.5; (b) stochastic gradient descent, learning rate 0.2; (c) AdaMax, learning rate 0.05; (d) noisy stochastic gradient descent, learning rate 0.2. Model parameters are $J_{1}=0.2$, $J_{2}=0.1$, $\Gamma=0.1$, $\gamma_{l}=0.1$.
@@ Appendix B Optimal energy landscapes: $N=5$ and $N=6$-site chains
Figure 14: Converged optimal energy profiles and corresponding population fluxes found when optimizing transport without environmental interactions. Parameters are the same as in Fig. 13.
@@
Fig. 16 shows that that the optimizer consistently converges to the same solution when starting from different initial conditions, with only mild variations across runs.
@@
Figure 16: Converged optimal energy profiles and corresponding population fluxes found when optimizing transport in OQS Model I. Parameters are the same as in Fig. 15.
@@
Figure 18: Converged optimal energy profiles and corresponding population fluxes found when optimizing transport in OQS Model II. Parameters are the same as in Fig. 17.

### v2

@@ IV.3 OQS Model II
We extend our investigation to transport optimization at finite temperatures using OQS Model II, as defined in Eq. (6). The results are shown in Fig. 10(a)–(b). Unlike the Coherent Model and OQS Model I, we observe consistent trends across both short- and long-range tunneling chains. In each case, the optimal energy profile adopts a ramp-like, steadily decreasing form. This configuration is physically intuitive: at moderate temperatures, it facilitates forward transfer from the entrance to the exit site while effectively suppressing backflow. For short-range tunneling, we also observe the emergence of high-energy levels above both the entrance and exit sites. Similarly to the three-site models shown in Figs. 4(c) and (f), this arrangement reduces population buildup along the chain itself while enhancing accumulation at the exit site.
Figure 10(c)-(d) show the population distributions corresponding to the optimized energy profiles, while Fig. 10(e)-(h) presents the off-diagonal elements of the density matrix. Interestingly, after the first one or two sites, the population in the chain increases towards the exit site, indicating on the accumulation of carriers towards the end of the chain due to the decreasing ramp potential. Compared to optimal structures identified in OQS Model I, coherences in the finite-temperature model are more localized, extending only to about three neighboring sites for both short- and long-range tunneling cases. To highlight the robustness of these findings, Fig. 11 compiles a set of energy profiles achieved during the optimization process, all consistently exhibiting the characteristic ramp-like structure.
@@
Model Figures Energy landscape Population profile Coherences
Coherent model, $\alpha=3$ 2, 5-6, 17-18 nearly flat nearly constant localized
Coherent model, $\alpha=1$ 2, 5-6, 17-18 corrugated bowl shaped extended
OQS Model I, $\alpha=3$ 3, 7-9, 19-20 nearly flat decaying approx. linearly localized
OQS Model I, $\alpha=1$ 3, 7-9, 19-20 corrugated constant for $n>4$ extended
OQS Model II, $\alpha=3$ 4,10-11, 21-22 up then down ramp growing towards exit site localized
OQS Model II, $\alpha=1$ 4, 10-11, 21-22 down ramp approx. constant at center localized
Table 1: Summary of models, energy landscapes optimizing transport, and the associated proposed transport mechanism, deduced based on the behavior of coherences and the population profile. “Localized” coherences refer to cases where quantum coherences are strongest between nearest-neighbor sites and decay rapidly with distance. In contrast, “extended” coherences describe situations where coherences with more distant sites can be comparable to or even stronger than those between immediate neighbors.
Figure 12: OQS Model I: Same as Fig. 7 but with a stronger dephasing constant, $\Gamma=0.2$. Other parameters are $J_{max}=0.2$ and $\gamma_{l}=0.1$.
Figure 13: OQS Model I: Same as Fig. 7 but with a smaller leakage rate, $\gamma_{l}=0.05$. Other parameters are $J_{max}=0.2$ and $\Gamma=0.1$.
IV.4 Discussion
We highlight that all optimization results were obtained under nonequilibrium steady state conditions and with a fixed value for the (physical) extraction rate $\gamma_{l}$. Our conclusions were drawn using a value for $\gamma_{l}$ comparable to other energy parameters. Having a very small $\gamma_{l}$ would turn it into the rate-determining step, while a very large $\gamma_{l}$ would effectively lead to a large level broadening at the exit site, shadowing variation in energy profile.
For OQS Model I, our analysis thus far has been performed using a fixed value of the environmental dephasing rate constant, $\Gamma$. It is important, however, to examine other values of $\Gamma$ to confirm the robustness of our observations. In Fig. 12, we present additional results for OQS Model I, analogous to those in Fig. 7, but obtained under stronger environmental noise. Overall, we find that the qualitative trends in the optimized energy profiles are preserved for both short-range and long-range tunneling systems. Interestingly, in the long-range tunneling case, the population profile at higher $\Gamma$ values shows a slight deviation from the flat behavior of Fig. 7, developing a small gradient across the bulk of the chain. This behavior aligns with the expectation that, as environmental effects become stronger relative to the tunneling amplitudes, long-range couplings, and their facilitation of coherent long-range hops, play a reduced role, giving way to more incoherent dynamics. We also confirmed convergence of the structure reported in Fig. 12 for many initial values, as in Fig. 8, and that the long-range system supports an ENAQT regime at $\alpha=1$, similarly to Fig. 9.
This example reinforces the rough classification of three transport regimes for the system, as discussed in Sec. II.5: coherent, incoherent and intermediate. Our choice of parameters throughout this project has been to achieve the intermediate regime, thus allowing nontrivial competition and optimization. The comparison of Fig. 12 with Fig. 7 clearly shows that increasing the dephasing effects relative to coherent effects takes us further from the intermediate regime to the incoherent limit.
We further examine the impact of changing $\gamma_{l}$ on the OQS Model I results. In Fig. 13, we use $\gamma_{l}=0.05$ compared to $\gamma_{l}=0.1$ in Fig. 7. We find that modification of the leak rate does not fundamentally change the results but that a smaller $\gamma_{l}$ leads to the accumulation of population near the exit site, which is expected. In contrast, we found that an increase in $\gamma_{l}$ (not shown) led to the depletion of the population at the exit site, compared to the bulk.
Overall, Figs. 7, 12 and 13 demonstrate the robustness of our observations when environmental noise or leak rate are varied, yet kept within the range of other parameters. Particularly we note that the optimization results are more robust when interactions are short-range.
@@ V Conclusions
Using modern optimization algorithms, we have identified distinct classes of energy landscapes that maximize population flux along quasi-one-dimensional chains. Hand-in-hand with the optimization, the analysis of the resulting steady state has allowed us to guess at underlying transport mechanisms. Toward the objective of optimized transport, we specifically studied (i) the role of the environment and (ii) the impact of short-range vs. long-range tunneling on chain design. The interplay of these incoherent and coherent effects leads to complex transport behavior. Optimized transport in the models required these factors to cooperate. Importantly, in the presence of long-range tunneling couplings, the optimization approach discovered energetic profiles that would be very difficult to guess analytically.
@@
We modeled the chain’s interaction with its environment using the Lindblad QME. In OQS Model I, we incorporated local Lindblad dephasing, which, when viewed in the energy eigenbasis, corresponds to coupling with an infinite-temperature bath. In contrast, OQS Model II represents a finite-temperature environment that drives both excitation and relaxation processes between energy eigenstates, while satisfying the detailed balance condition.
@@
We have identified the following design rules for optimizing transfer, summarized in Table I.
@@
(i) Coherent model: for short-range tunneling, the optimal configuration features an almost uniform energy landscape, enabling a constant population profile and thus near-resonant transport. When long-range tunneling contributes, the optimal energy profile becomes highly non-uniform and corrugated, with intermediate sites energetically detuned. This facilitates a more direct transfer between the chain’s endpoints, as evidenced by strong steady-state coherences linking those sites. Long-range coherent tunneling systems are highly sensitive to variations in parameters near optimal energy landscapes.
@@
(ii) OQS Model I: when local dephasing is comparable in magnitude to the tunneling coupling energies, models with short-range tunneling favor nearly flat energy landscapes, with only slight detuning from the entry and exit sites to facilitate conduction. The population profile in this case suggests diffusive transport. In contrast, under long-range tunneling, a corrugated energy landscape allows the system to cross over to quasi-ballistic conduction away from the entrance site by suppressing interferences and facilitating efficient end-to-end transfer. The development of this motion under highly corrugated structures and with dephasing effects manifests that the optimization approach can discover patterns that would otherwise be impossible to determine analytically.
@@
(iii) OQS Model II: at moderate temperatures ($T\approx J_{max}$), ramp-like energy profiles with a total gap of the order of $J_{max}$ are preferred, as they support directional transport and improve the collection of carriers at the exit site. In these systems, regardless of the range of the tunneling couplings, transport proceeds in an almost localized manner with small coherences, and with population concentrating towards the bottom (end) of the ramp.
We emphasize that the transport mechanisms discussed in this work are inferred only from the observed patterns of coherences and population profiles. Because the energy landscape was optimized separately for structures of different lengths, potentially leading to distinct profiles for different sizes, we cannot systematically examine how the flux scales with system size for identical Hamiltonians. Thus, a quantitative analysis of transport mechanisms is beyond the scope of this study. Our objective, instead, has been to identify families of systems that enhance flux under different conditions, rather than to classify transport mechanisms in detail.
@@
The optimization process posed challenges at times, especially for the Coherent Model and OQS Model II. In the coherent model, particularly with long-range tunneling, the steady-state populations were highly sensitive to small changes in energy levels, resulting in complex, rugged energy landscapes with numerous local maxima that complicated optimization. For OQS Model II, the need to re-diagonalize the system Hamiltonian at each energy iteration further increased computational demands, making the optimization considerably slower compared to OQS Model I under the same parameters.
@@
The concept of ENAQT was originally introduced to explain the efficiency of light-harvesting systems [75, 79, 12, 16]. It was later generalized to broader scenarios of transport in low-dimensional chains [70, 8]. We believe that the main impact of our work lies in providing a deeper understanding of transport in such engineered setups, with potential relevance for applications such as photovoltaics. Long-range steady-state coherences, such as those found under OQS Model I in Fig. 7(f)-(h) have been shown to lead to robust entanglement [102, 32, 56]. As such, the energy profiles we obtain under long-range tunneling [purple in Fig. 7(a)-(b)] may find applications in quantum information processing and quantum metrology [32, 96, 55].
@@
Future work could extend this study of carrier transfer to multi-carrier scenarios, such as transport in spin systems modeled by Heisenberg-type chains with long-range exchange interactions. Another promising avenue is the exploration of excitation transfer in molecular aggregates or disordered materials coupled to optical cavities [71, 92, 36, 113, 112], where cavity interactions mediate long-range transfer. Joint optimization of both the molecular structure and its cavity coupling may uncover new design principles for efficient carrier transport in quantum networks.
In sum, this work establishes a framework for designing quantum systems that leverage the synergy of coherent and incoherent processes to optimize transport performance. Our results provide guiding principles for engineering energy profiles to achieve efficient transport, with potential applications in organic photovoltaics, nanoelectronics, and quantum networks.
@@ Acknowledgements.
Data Availability. The data that support the findings of this article are openly available [64]. The underlying code will be made available early in 2026.
@@ Appendix A Details of the optimization process
Stochastic Gradient Descent (SGD) and Noisy SGD algorithms were also tested with three sites, the results of which are shown in Figs. 14(b) and 14(d), respectively. Both found the same optima as OGA and AdaMax, but with more steps: SGD took between 890 and 17,700 steps, and with the default noise levels, Noisy SGD took between 19,000 and 28,000 steps.
We emphasize that our work does not employ, nor does it develop, machine learning models. Instead, we use gradient-based optimization methods, specifically, gradient ascent, whose efficiency is enhanced through automatic differentiation (AD). Although AD was developed alongside modern machine learning, it is a general computational technique for exactly and efficiently propagating derivatives through general programs via the chain rule. It enables rapid and accurate evaluation of gradients, therefore automating and accelerating numerical optimization. However, it does not involve statistical learning or model training.
@@
Figure 14: Examples of optimization trajectories in a three-site system under OQS Model I. Squares mark the initial energies supplied to the optimizer, while stars indicate where the optimizers reached their stopping condition. Default Optax parameters were used for each algorithm, except for the learning rate. The optimizers are: (a) optimistic gradient ascent, learning rate 0.5; (b) stochastic gradient descent, learning rate 0.2; (c) AdaMax, learning rate 0.05; (d) noisy stochastic gradient descent, learning rate 0.2. Model parameters are $J_{1}=0.2$, $J_{2}=0.1$, $\Gamma=0.1$, $\gamma_{l}=0.1$.
@@ A.3 OQS Model II
Sharpness of Maxima Optimizing the system energies under the Coherent Model (CM) of transport was difficult. These optimizations took more steps, and fewer initial energy configurations successfully converged to local maxima. To confirm our optimal profile was indeed a maximum, we study the effects of varying $\varepsilon_{4}$ in the best 10-site, long-range tunneling system optimized under the CM while leaving the other energies fixed at their optimal values. This is shown in Figure 15. The maximal population flux does indeed occur at the value of $\varepsilon_{4}$ found by the Adamax algorithm, though this peak is very narrow, with a steep slope to its left. There are other local maxima near $\varepsilon_{4}\approx 0$, though they are also sharply peaked. In contrast, when we perform the same variation in $\varepsilon_{4}$ with the best 10-site, long-range tunneling system optimized under OQS I, shown in Figure 16, we see a much smoother maximum and no other local maxima in the range of energies considered.
Figure 15: (a) Coherent Model: the optimized energy landscape shown in Fig. 5(b) for long-range ($\alpha=1$) tunneling. The black bar shows the range over which $\varepsilon_{4}$ was varied. (b) Population flux $\eta$ as a function of $\varepsilon_{4}$. The dashed line shows the value of $\varepsilon_{4}$ found through the optimization procedure. Other parameters are $J_{max}=0.2$ and $\gamma_{l}=0.1$.
Figure 16: (a) OQS Model I: the optimized energy landscape shown in Fig. 7(b) for long-range ($\alpha=1$) tunneling. The black bar shows the range over which $\varepsilon_{4}$ was varied. (b) Population flux $\eta$ as a function of $\varepsilon_{4}$. The dashed line shows the value of $\varepsilon_{4}$ found through the optimization procedure. Other parameters are $J_{max}=0.2$, $\Gamma=0.1$, and $\gamma_{l}=0.1$.
@@ Appendix B Optimal energy landscapes: $N=5$ and $N=6$-site chains
Figure 18: Converged optimal energy profiles and corresponding population fluxes found when optimizing transport without environmental interactions. Each column in any panel is a single profile, and the leftmost columns are those plotted in Fig. 17(a)-(b). Parameters are the same as in Fig. 17.
@@
Fig. 20 shows that the optimizer consistently converges to the same solution when starting from different initial conditions, with only mild variations across runs.
@@
Figure 20: Converged optimal energy profiles and corresponding population fluxes found when optimizing transport in OQS Model I. Each column in any panel is a single profile, and the leftmost columns are those plotted in Fig. 19(a)-(b). Parameters are the same as in Fig. 19.
@@
Figure 22: Converged optimal energy profiles and corresponding population fluxes found when optimizing transport in OQS Model II. Each column in any panel is a single profile, and the leftmost columns are those plotted in Fig. 17(a)-(b). Parameters are the same as in Fig. 21.